OpenStudy (anonymous):

Differentiate f(x)=yln(6+e^y)

OpenStudy (jamesj):

f(x)=yln(6+e^y), not x? but y?

OpenStudy (anonymous):

yea... this is one of my issues

OpenStudy (jamesj):

It's a trick. With no other information, then y here is just a constant and hence f(x) it itself a constant function. Hence df/dx = 0.

OpenStudy (anonymous):

not correct. just input it into webassign.

OpenStudy (jamesj):

In which case either it's an error and you should consider all the ys to be xs. Or, you should consider y as a function of x itself and use the chain rule/implicit differentiation. I'd try the first of those two first.

OpenStudy (anonymous):

my bad its f(y)=....

OpenStudy (jamesj):

there you are then.

OpenStudy (anonymous):

so what do i need to do it it? I've only got like one more try. and I dont come here until its almost too late lol. try to swing it on my own

OpenStudy (jamesj):

Chain rule and product rule. You need to try and sweat this out; it's how you learn.

OpenStudy (anonymous):

i'm working on it lmao! do you go to college still? whats your major?

OpenStudy (jamesj):

For instance let u = y, v = ln(6+e^y) Then df/dy = du/dy . v + u . dv/dy Now you need to figure out how to differentiate v, and that's the chain rule.

OpenStudy (jamesj):

I'm a mathematician slumming it.

OpenStudy (anonymous):

i'm majoring in math right now... want to do applied and industrial math. Hows the pay? and how hard is the job market?

OpenStudy (jamesj):

Impossible questions to answer in general, but in short good mathematicians are in short supply; and good ones can do this problem in their heads. go back to this question you asked: http://openstudy.com/groups/mathematics#/groups/mathematics/updates/4e9a31410b8ba77642a9c5ee Make sure you understand every step. If you do, you should have few problems here.

OpenStudy (anonymous):

well i got my ged in 1997 and started college in 2009 after a few times of jail and stupid stuff. so in my head aint quite there yet. but it comes decently easy once shown in detail once. Thanks for your time!

OpenStudy (anonymous):

so i got 1+ 1/(6+e^y)*e^y and that still aint right. guess i'll just bring it to my professor

myininaya (myininaya):

is suppose to say f(y)=yln(6+e^y) ?

myininaya (myininaya):

\[f'(y)=(y)'\ln(6+e^y) +y[\ln(6+e^y)]'\]

OpenStudy (anonymous):

I did it !!!!!!!!!!!!!!!!!! thanks myininaya and JamesJ

myininaya (myininaya):

np